OpenStudy (bugman):

how do I change this equation from standard form into vertex and intercept form? y=x^2-4x-5

OpenStudy (joshmorris):

Hey

OpenStudy (bugman):

hey, Can you help? please

OpenStudy (joshmorris):

y = x^2 - 4x - 5 = x^2 - 4x + 4 - 4 - 5 = (x^2 - 4x + 4) -4 - 5 y = (x-2)^2 - 9 Here, the vertex is at (2, -9)

OpenStudy (joshmorris):

x = 2 (the vertical line that goes through the vertex)

OpenStudy (joshmorris):

The graph of this function is called a parabola. To graph a parabola, use its vertex (in this case, (2, -9) ) and y-intercept (y = -5). The y-intercept is the constant in the given equation. Use the fact that the vertical line through the vertex is a line of symmetry. Plot the point corresponding to the y-intercept that lies on the other side of the vertical line x = 2.

OpenStudy (joshmorris):

The graph is translated 2 units to the right and 9 units down.

OpenStudy (bugman):

so then y=(x-2)^2-9 is vertex form, and a parabola is the curve line I'm pretty sure

OpenStudy (joshmorris):

Yes

OpenStudy (joshmorris):

Medal?

OpenStudy (bugman):

ok, great now how do I do the intercept form?